OpenStudy (anonymous):

The center of a circle is (4, 6), and an endpoint of a diameter is (2, 5). What is the other endpoint of the diameter? is it (3, 5.5)? not sure..

jimthompson5910 (jim_thompson5910):

Hint: let the other point be (x,y) This would mean (x+2)/2 = 4 and (y+5)/2 = 6

jimthompson5910 (jim_thompson5910):

What do you get when you solve (x+2)/2 = 4 for x?

OpenStudy (anonymous):

(6, 7)

jimthompson5910 (jim_thompson5910):

better, but hopefully you're not guessing

jimthompson5910 (jim_thompson5910):

and you can see how you got that answer